OpenStudy (anonymous):

which ordered pair is a solution of the equation 3x + y = -23? A. (5, 8) B. (-8, -5) C. (-5, -8) D. (8, 5)

OpenStudy (whpalmer4):

Which ordered pair makes the equation work? Let's try D: \[x=8,y=5\]\[3x+y=-23\]\[3(8)+5=-23\]\[24+5=-23\]\[29=-23\]Nope, it isn't D!

OpenStudy (whpalmer4):

Hopefully it is obvious from that experiment that you'll need values of \(x\) and \(y\) that are negative...

OpenStudy (anonymous):

B?

OpenStudy (whpalmer4):

well, does it work? \[x=-8,y=-5\]\[3(-8)+(-5) = -23\]Is that true or not?

OpenStudy (anonymous):

umm, i'm not sure

OpenStudy (whpalmer4):

what is 3(-8)?

OpenStudy (anonymous):

-24

OpenStudy (whpalmer4):

right. what is -24 - 5?

OpenStudy (anonymous):

-29

OpenStudy (whpalmer4):

right. does -29 = -23?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

okay. is B a correct answer?

OpenStudy (anonymous):

so it is A or C?

OpenStudy (whpalmer4):

go through the process. find the values of x and y that make the equation true. there's no need to guess, just do a little arithmetic.